OpenStudy (anonymous):

Find the value of x so that the line passing through (x, 10) and (-4, 8) has a slope of

OpenStudy (misty1212):

oops

OpenStudy (anonymous):

2/3

OpenStudy (anonymous):

@misty1212 my bad!

OpenStudy (misty1212):

ok ok now now we can get somewhere the slope is found by computing \[\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

how would i do hat without x

OpenStudy (misty1212):

\[\frac{10-8}{x+4}=\frac{2}{3}\]

OpenStudy (anonymous):

wait how'd you get 3 from the 4+x?

OpenStudy (misty1212):

you gotta find the \(x\) but you can probably figure it out without using algebra

OpenStudy (anonymous):

can u explain

OpenStudy (misty1212):

you said the slope was \(\frac{2}{3}\) right?

OpenStudy (anonymous):

oh right

OpenStudy (misty1212):

so i computed the slope using the formula and put it equal to 2/3 the only number we do not know is x

OpenStudy (misty1212):

i hope that explains this line \[\frac{10-8}{x+4}=\frac{2}{3}\] you still need to find x

OpenStudy (misty1212):

but it is now pretty easy \[\frac{10-8}{x+4}=\frac{2}{3}\\ \frac{2}{x+4}=\frac{2}{3}\] so \[x+4=3\] and then you find x

OpenStudy (anonymous):

the 2nd step to the third kinda threw me off

OpenStudy (misty1212):

\[\frac{2}{x+4}=\frac{2}{3}\\ x+4=3\]? that part?

OpenStudy (misty1212):

ooh look at all the ghosts, must be halloween

OpenStudy (danjs):

why does the site do that?

OpenStudy (anonymous):

i was wondering what is that???

OpenStudy (misty1212):

idk

OpenStudy (anonymous):

do u guys know what it is // why

OpenStudy (misty1212):

ghosts of openstudy past maybe

OpenStudy (danjs):

Thought it was my browser, guess not

OpenStudy (anonymous):

but about the math yes that part

OpenStudy (anonymous):

yeah same

OpenStudy (anonymous):

could it be the site checking up

OpenStudy (misty1212):

you are thinking like a confused math student not a person that is why (not to be rude or anything)

OpenStudy (misty1212):

\[\frac{2}{x+4}=\frac{2}{3}\] the top of both is 2 right?

OpenStudy (misty1212):

since the numerators are both 2, the only way for them to be equal is if the denominators are both 3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

sorry I'm stupid

OpenStudy (misty1212):

lol no not stupid, just so stuck on the other stuff that you miss the totally obvious

OpenStudy (misty1212):

math class has a way of doing that to otherwise completely intelligent people

OpenStudy (anonymous):

very true

OpenStudy (anonymous):

what would be the value of x? @misty1212

OpenStudy (misty1212):

\[x+4=3\\ x=?\]

OpenStudy (anonymous):

-1

OpenStudy (misty1212):

bingo

OpenStudy (anonymous):

thank u for all your help @misty1212

OpenStudy (misty1212):

you are welcome dear any time